OpenStudy (thatawayz):

Geometry: Basic Trig Help Please, Ill list question, i dont want answer i want help to solve it.

OpenStudy (thatawayz):

(07.04 MC) An observer (O) spots a plane flying at a 35° angle to his horizontal line of sight. If the plane is flying at an altitude of 17,000 ft., what is the distance (x) from the plane (P) to the observer (O)?

OpenStudy (thatawayz):

Hight is 17,000 ft, angle is 35 degrees, we are trying to find the hypotenuse

OpenStudy (anonymous):

First step is to draw it out, so do that.

OpenStudy (thatawayz):

Alright, hold up

OpenStudy (thatawayz):

Okay, drawn out

OpenStudy (anonymous):

Ok, good. So relative to the angle, is the altitude side opposite or adjacent?

OpenStudy (thatawayz):

the altitude side would be oppisite

OpenStudy (anonymous):

And what trig function has the opposite and the hypotenuse in it?

OpenStudy (thatawayz):

Sine, Right?

OpenStudy (anonymous):

Right. So if O=17,000 and H, how would we construct our formula?

OpenStudy (anonymous):

Shoulda said "and we don't know H"

OpenStudy (thatawayz):

Wouldnt it be Sine Of 35 degrees is equal to H "which in this case is 17,000", over x "the lenght of the hypoteneuse

OpenStudy (anonymous):

Did you mean to put an O there instead of an H to represent the opposite side? Cause if you did, that looks fine. In equation form though, what would that be?

OpenStudy (thatawayz):

By the way, thanks for helping me and im sorry if this takes a while, i am not good with mathematics

OpenStudy (anonymous):

No problem, we're moving pretty quick. I spent over an hour helping someone with probability earlier.

OpenStudy (thatawayz):

Ah, i see, well, lets hope i dont take too long, okay, did i right it down correctly?

OpenStudy (anonymous):

Well right, but I was looking for it in more of an equation form. Like this \[\large \sf sin(35)=\frac{17000}{H}\]

OpenStudy (thatawayz):

ohh, im sorry, i did not know you can write out the equation like that

OpenStudy (anonymous):

Well anyway, do you know how to solve for H from here?

OpenStudy (thatawayz):

you multiply by both sides and it cancels out. so you get\[h \times \sin35=17000/h\]

OpenStudy (anonymous):

Well not really. When you have \[\large \sf a=\frac{b}{c}\] and you want to get c alone, just switch a and c \[\large \sf c=\frac{b}{a}\] So for \[\large \sf sin(35)=\frac{17000}{H}\] it would be \[\large \sf H=\frac{17000}{sin(35)}\]

OpenStudy (thatawayz):

So flip it like a reciprocal?

OpenStudy (anonymous):

You're just switching a and c, to get c alone. In this case, we're getting H alone, like I just showed

OpenStudy (thatawayz):

Alright, hold up, im writing this down

OpenStudy (thatawayz):

okay i wrote it down, so H equals 17000 over sin 35

OpenStudy (thatawayz):

flip a and c

OpenStudy (anonymous):

Yup, so just plug that into a calculator and you'll have your answer

OpenStudy (thatawayz):

alright, now this is for the hypoteneuse correct? cause thats what im trying to find

OpenStudy (anonymous):

Yes, the H represents the hypotenuse

OpenStudy (thatawayz):

alright, thanks for the clarification

OpenStudy (anonymous):

No problem, just tell me what you get when you get it

OpenStudy (thatawayz):

Okay, im really sorry, how would i plug this in a calculator, in the same format given?

OpenStudy (anonymous):

Just do \[\large \sf \frac{17000}{sin(35)}\]

OpenStudy (thatawayz):

and press the execute button on the calculator?

OpenStudy (anonymous):

if that's the button that executes it, then yes

OpenStudy (thatawayz):

ok

OpenStudy (thatawayz):

okay, i got-39702.68115

OpenStudy (thatawayz):

which i am almost certain is not being displayed correctly, perhaps user error on my part

OpenStudy (anonymous):

Then your calculator is in radians, I forgot to tell you it must be in degrees. My apologies.

OpenStudy (thatawayz):

no worries, what needs to be in degrees exactly, the 35?

OpenStudy (anonymous):

The setting of the calculator.

OpenStudy (thatawayz):

ok sin of 35 degrees is roughly 0.57

OpenStudy (thatawayz):

the answer is 29824

OpenStudy (anonymous):

Good, although specifically it is 29,639. But I think you got the point of it

OpenStudy (thatawayz):

alright i see the point of this one, thankyou for taking the time to help me, n ow im new to this site so is there any way of giving kudos to someone who has helped?

OpenStudy (anonymous):

You can press the "Best Response" button for me, which will give me a medal. (:

OpenStudy (thatawayz):

alright thanks again, i do have more questions but i dont want to take anymore of your time

OpenStudy (anonymous):

And if you feel you may need future help you can fan me so you can tag me in the future. You could fan zepdrix too, he's real smart

OpenStudy (anonymous):

Oh and no problem, have a nice rest of your day (:

OpenStudy (thatawayz):

alright thanks, i have aleready became a fan of you

OpenStudy (thatawayz):

Closing question